OpenStudy (anonymous):

Logan's only debt obligations are a car loan payment of $512 and a credit card payment of $70 every month. What is the minimum amount of money he must take home every month in order to avoid being in danger of credit overload? A. $2328 B. $2560 C. $582 D. $2910

OpenStudy (tkhunny):

You will need a rule for that. 30%? 40%?

OpenStudy (tkhunny):

512+70 = 582 582/2328 = 0.25 582/2560 = 0.22734375 582/582 = 1.00 582/2910 = 0.2 If your rule is anything but 20%, 22.734%, 25%, or 100%, you are out of luck.

OpenStudy (anonymous):

Do you know that rule?

OpenStudy (anonymous):

yikes!

OpenStudy (anonymous):

it was 2910